# Benchmarking LLM Privacy Recognition for Social Robot Decision Making

### v1

@@ Abstract.
Social robots are embodied agents that interact with people while following human communication norms. These robots interact using verbal and non-verbal cues, and share the physical environments of people. While social robots have previously utilized rule-based systems or probabilistic models for user interaction, the rapid evolution of large language models (LLMs) presents new opportunities to develop LLM-empowered social robots for enhanced human-robot interaction. To fully realize these capabilities, however, robots need to collect data such as audio, fine-grained images, video, and locations. As a result, LLMs often process sensitive personal information, particularly within home environments. Given the tension between utility and privacy risks, evaluating how current LLMs manage sensitive data is critical. Specifically, we aim to explore the extent to which out-of-the-box LLMs are privacy-aware in the context of household social robots. In this study, we present a set of privacy-relevant scenarios crafted through the lens of Contextual Integrity (CI). We first survey users’ privacy preferences regarding in-home social robot behaviors and then examine how their privacy orientation affects their choices of these behaviors ($N=450$). We then provide the same set of scenarios and questions to state-of-the-art LLMs ($N=10$) and find that the agreement between humans and LLMs is low. To further investigate the capabilities of LLMs as a potential privacy controller, we implement four additional prompting strategies and compare their results. Finally, we discuss the implications and potential of AI privacy awareness in human-robot interaction.
@@ 1. Introduction
As robots become more competent and provide greater utility, we increasingly rely on them to perform tasks in private locations such as homes (New York State Office for the Aging, 2025) and care settings  (Jack, 2024; The Sun Tech, 2025). As early as 2014, more than 3,000 Paro therapeutic robots had been sold worldwide (National Institute of Advanced Industrial Science and Technology (2013), AIST), and by 2024, over 50 million Roomba units had been sold globally (iRobot Corporation, 2025). At present, the market size for social robots has increased to \$5.6 billion, with a forecast of \$42.5 billion in 2033 (IMARC Group, 2024).
Simultaneously, researchers in academia and industry are implementing large language models (LLMs) and video language models (VLMs) on robots. These models equip robots with the tools to perceive and reason about their environments far beyond their previous capabilities. In recent years, researchers have demonstrated that LLMs can enhance robot capabilities by enabling natural language requests (Ahn et al., 2022), writing code to control robot movement on-the-fly (Liang et al., 2023), achieving interactive dialogue (Shinn et al., 2023), and improving navigation abilities (Hassan et al., 2024). Additionally, researchers have evaluated LLM-powered robots through user studies to identify the potential benefits and weaknesses of these LLM frameworks (Yuan et al., 2024a; Murali et al., 2023; Kim et al., 2024a; Azeem et al., 2024). Several large tech companies have also introduced projects that utilize LLMs or VLMs to enable more intelligent robots, such as Google’s DeepMind-RT-2 (Brohan et al., 2023) and NVIDIA’s GR00T N1 (Bjorck et al., 2025). From these advancements, it is clear that LLMs and VLMs can offer significant benefits when used in conjunction with robots.
@@
Like many other technologies, robots pose a threat to users’ data privacy. Robots can autonomously navigate human environments and collect data in private spaces that users may not have originally intended or desired. When a robot captures audio or visual data of humans or human environments, that data may be collected by the robot’s manufacturers. While this data collection may amount to a privacy violation on its own, further use in training, marketing, or leakage of data can further cause harm to the individuals from which the data originated. Beyond these traditional data privacy concerns, however, social robots pose an additional threat to users’ interpersonal privacy. Social robots can interact with users in a human-like manner that encourages engagement and communication. Once a robot has captured user data in this interpersonal context, it then has the potential to share that data with other users. This phenomenon is unique to social robots and places them in a position to engage in privacy decision-making. It is within this unique context that LLMs may have the potential to alleviate privacy concerns. LLMs may be able to determine whether a given piece of information is private, whether it is necessary to retain (e.g., for functional purposes), and whether or with whom it may be shared. Privacy is highly contextually dependent, and the nuances of any given scenario may require these highly sophisticated models to navigate appropriately.
@@
Though many state-of-the-art LLMs have been designed with some level of privacy protection in mind, it is not yet clear the degree to which these models are privacy aware. Prior works have primarily focused on general LLM safety, such as LLMs’ vulnerability against jailbreak attacks (Zhuo et al., 2023). Few works, however, have investigated the privacy perceptions of LLMs and VLMs (Mireshghallah et al., 2023; Shao et al., 2024), and none have done so in the context of human-robot interaction. Given this gap, we aim to investigate whether and how LLMs interpret privacy scenarios in the context of human-robot interaction in home environments. Specifically, we raise the following research questions: RQ1: How does an individual’s privacy orientation influence their privacy expectations of social robots?; RQ2: How well do state-of-the-art LLMs align with individuals’ privacy expectations?; and RQ3: How do prompting strategies influence state-of-the-art LLMs’ conformity with individuals’ privacy expectations? Through the development and evaluation of privacy scenarios with participants and LLMs, we show that state-of-the-art LLMs maintain a broad understanding of privacy sensitivity, but may not yet possess nuanced privacy awareness.
In addressing our research questions, we present the following contributions:
  • Data Set: A set of privacy scenarios grounded in contextual integrity (CI)
  • Annotation: Multiple large online user studies to determine appropriate robot and LLM responses to scenarios
  • Evaluation: A benchmark of 10 state-of-the-art LLMs’ privacy awareness using multiple prompting strategies
  • Design Implications: Insights into the use of LLMs as potential privacy-controllers for social robots
@@ 2. Related Works
We survey prior work on the intersection of LLMs, privacy, and robotics to present an overview of current research directions relevant to our work. As we aim to explore whether user privacy preferences align with LLM-powered robots, we introduce the following threads of existing literature: (1) conceptual approaches to privacy, including privacy orientations and contextual integrity, (2) privacy concerns related to robot data collection in homes, (3) the integration of LLM reasoning in robots, and (4) LLM privacy risks and benchmarking.
@@ 2.1. Conceptual Approaches to Privacy
Beyond this broad framework of privacy, many other works have attempted to assess the specific privacy attitudes, orientations, and personas of individuals as they engage with technology (Preibusch, 2013). These efforts have focused on online privacy (Elueze and Quan-Haase, 2018; Buchanan et al., 2007), out-of-device privacy (i.e., privacy violations through direct human observation) (Farzand et al., 2024), privacy perceptions broadly (Chignell et al., 2003; Giang et al., 2023), and the identification of privacy personas (Hrynenko and Cavallaro, 2024). Of particular note to this work is the Privacy Orientation Scale (POS) (Baruh and Cemalcılar, 2014), which captures an individual’s perceptions of privacy as a right, concerns about their own informational privacy, other-contingent privacy, and concerns about the privacy of others. These four subscales are then used to assess a user’s privacy orientation, including privacy advocates, privacy individualists, and privacy indifferents. While our results do not fit neatly within these privacy orientations, as we will see in Section 4.4.1, other works have described the general challenge of defining privacy personas (Biselli et al., 2022). Given the substantial contextual dependence of privacy, domain-agnostic assessments of privacy perceptions may also struggle to generate consistent personas across domains.
@@ 2.2. Privacy in Robotics
While privacy has been explored in a myriad of fields, it is a growing area of concern in robotics (Denning et al., 2009; Guerrero, 2017; Rueben et al., 2018; Torresen, 2018; Pagallo et al., 2018; Lutz et al., 2019; Chatzimichali et al., 2020). Particular emphasis has been placed on robotics usage in homes (Levinson et al., 2024b, a; Fernandes et al., 2016; Urquhart et al., 2019), as these environments are highly private, contain substantial amounts of sensitive information, and often offer inhabitants a sense of security. Given the capacity of some robots to capture data, autonomously navigate environments, and socially engage with humans, they pose a significant threat to privacy with unique challenges compared to many other technologies. Existing works have also focused on the inherent privacy-utility trade-off that exists when robots are introduced into new environments (Butler et al., 2015; Leite and Lehman, 2016). While users may have concerns about robot privacy violations, they may be tolerated if sufficient utility is provided by the robot (Lutz and Tamó-Larrieux, 2020).
Though these concerns are widespread, many prior works have proposed solutions that attempt to limit data collection, retention, and exposure (Sullivan and Mutlu, 2025). For example, low-resolution sensors (Kim et al., 2019) and purposeful sensor selection (Eick and Antón, 2020) have been proposed to limit unnecessary data collection, while sensor redirection (Shome et al., 2023) and privacy controllers (Tang et al., 2022) have been proposed to mitigate data exposure. While these efforts have uncovered new methods for preserving human privacy, the combination of LLMs and robots presents novel opportunities for privacy-preserving LLM decision-making.
@@ 2.3. Robots & LLM Integration
While mainstream LLMs are a relatively recent development (Vaswani et al., 2017; Radford et al., 2018), LLM integration in robots has been a clear area of interest for HRI applications. Though this intersection of technologies is still in its infancy, several works have already utilized LLMs in robotics applications. These applications have been specialized, such as a chemistry assistant robot (Yoshikawa et al., 2023) and a health attendant robot (Kim et al., 2024b), as well as broad, such as a ‘robotic brain’ to manage memory and control (Mai et al., 2023). Other works have considered the potential of LLM-powered robots compared to other forms of LLMs (i.e., text- and voice-based) (Kim et al., 2024a). All of these works, however, only begin to explore the full potential of robot-LLM integration. Further study is needed to understand the implications of LLMs for perception, control, interaction, and, most relevant to this work, decision-making (Zeng et al., 2023).
@@ 2.4. LLM Privacy Risks & Benchmarking
To assess the potential risks and benefits that LLMs present, researchers utilize benchmarking as a standardized evaluation method for LLM performance. Common evaluation domains include reasoning (Srivastava et al., 2022; Hendrycks et al., 2020), multilingualism (Hu et al., 2020; Liang et al., 2020), coding (Chen et al., 2021; Austin et al., 2021), and instruction following (Zheng et al., 2023; Dubois et al., 2025) as well as safety and privacy (Ruan et al., 2023; Yuan et al., 2024b; Zhang et al., 2023; Gu et al., 2024; Zhang et al., 2024b; Zharmagambetov et al., 2025; Zhu et al., 2024; Zhang et al., 2024a; Chang et al., 2025). Additional benchmarks, such as PURE (Chi et al., 2023), PolicyQA (Ahmad et al., 2020), PrivacyQA (Ravichander et al., 2019), and GenAIPABench (Hamid et al., 2023), have been performed to evaluate privacy-policy comprehension using dedicated datasets. Of particular relevance to this work, Mireshghallah et al. (2023) first raised the question of whether LLMs can reason about CI. These authors utilized their benchmark, ConfAlde, to show that GPT-4 violates CI norms in 39% of evaluated scenarios where humans would withhold the information. Most recently, PrivacyLens (Shao et al., 2024) introduced vignette-based questions to survey LLMs’ privacy-norm awareness. In this work, we aim to investigate the privacy awareness of LLMs by evaluating them through a benchmark grounded in CI.
@@
Robot Involvement Robot enters part-way, Robot present from the beginning, Robot collects data from afar (i.e, spying and eavesdropping)
@@
Table 1. Dimensions of Contextual Privacy—Dimensions of privacy relevant to household HRI. These dimensions were originally derived from literature on Contextual Integrity and expanded to include subdimensions and factors that are of particular relevance to HRI.
@@ 3. Developing Privacy Scenarios
To address our research questions, we aimed to develop a set of privacy-relevant scenarios that allow participants to report their level of concern with a robot collecting user data and their preferences for how that robot should behave following data collection. Here we describe our process to develop these scenarios, including the use of two online user studies to crowdsource specific scenario characteristics and robot behaviors. The studies presented below were each approved by the institutional review board of the University of Wisconsin–Madison, and all participants provided consent prior to their participation.
@@ 3.1. Dimensions of Contextual Integrity
To begin, we first reviewed existing literature on general in-home privacy concerns (Choe et al., 2011) and Contextual Integrity (Nissenbaum, 2004). From this literature, we became oriented with a comprehensive list of the many in-home activities individuals wish to keep private. Additionally, we distilled the primary dimensions of contextual privacy, including the context, actors, attributes, transmission principles, and purposes. From each of these dimensions, we further defined subdimensions relevant to privacy in human-robot interaction. These expanded dimensions include type of location, robot involvement, number of users, age of users, relation of users, medium of data, protected status of data, information flow, and task dependence. Each of these dimensions contains between two and four factors, and a full listing of all original dimensions, expanded dimensions, and factors can be seen in Table 1.
While most of these dimensions and their factors are fairly intuitive, the factors of protected status and medium of data are worth discussing further. Our original intent behind protected status was to capture the sensitivity of the information involved in a given scenario. Given how subjective sensitivity may be, we chose to instead define protected status as either protected or non-protected based upon legal norms. For example, information we consider to be protected includes health data (for Civil Rights, 2003), academic records (U.S. Congress, 1974), and legal records (United States Courts, 2008), among others (Internal Revenue Service, 1976; U.S. Congress, 1999). Conversely, we consider information such as vacation plans or intimacy between partners to be non-protected. Additionally, we considered only specific mediums of data that we believed an LLM or VLM could reliably recognize. These data types include visuals of the human body in compromising situations (e.g., bathing, toileting, or intimacy), visuals of device screens, audio of conversations, and visuals of physical documents. These broad categories were selected for the wide variety of information each of them could convey to an observer.

### v2

@@ Abstract.
While robots have previously utilized rule-based systems or probabilistic models for user interaction, the rapid evolution of large language models (LLMs) presents new opportunities to develop LLM-powered robots for enhanced human-robot interaction (HRI). To fully realize these capabilities, however, robots need to collect data such as audio, fine-grained images, video, and locations. As a result, LLMs often process sensitive personal information, particularly within private environments, such as homes. Given the tension between utility and privacy risks, evaluating how current LLMs manage sensitive data is critical. Specifically, we aim to explore the extent to which out-of-the-box LLMs are privacy-aware in the context of household robots. In this work, we present a set of privacy-relevant scenarios developed using the Contextual Integrity (CI) framework. We first surveyed users’ privacy preferences regarding in-home robot behaviors and then examined how their privacy orientations affected their choices of these behaviors ($N=450$). We then provided the same set of scenarios and questions to state-of-the-art LLMs ($N=10$) and found that the agreement between humans and LLMs was generally low. To further investigate the capabilities of LLMs as potential privacy controllers, we implemented four additional prompting strategies and compared their results. We discuss the performance of the evaluated models as well as the implications and potential of AI privacy awareness in human-robot interaction.
@@ 1. Introduction
As robots become more competent and provide greater utility, we increasingly rely on them to perform tasks in private locations such as homes (New York State Office for the Aging, 2025) and care settings  (Jack, 2024; The Sun Tech, 2025). As early as 2014, more than 3,000 Paro therapeutic robots had been sold across 30 countries (of Advanced Industrial Science and, AIST), and by 2024, over 50 million Roomba units had been sold globally (iRobot Corporation, 2025). In 2024, the market size for consumer robots increased to \$10.92 billion, with a forecast of \$40.15 billion in 2030 (Grand View Research, 2024).
@@
Like many other technologies, robots pose a threat to users’ data privacy (Guo, 2022). Robots can autonomously navigate human environments and collect data in private spaces that users may not have originally intended or desired. Beyond traditional data privacy concerns that robots pose as a connected device collecting audio or visual data (Fell, 2024), they introduce a new threat to users’ interpersonal privacy (Tang et al., 2022; Bell et al., 2025). Robots can interact with users in a human-like manner, encouraging engagement and communication. Once a robot has captured user data in this interpersonal context, it can then potentially share that data with other users. This phenomenon is unique to robots in home settings, placing them in a critical position to engage in privacy decision making. It is within this unique context that large language models (LLMs) may have the potential to alleviate privacy concerns in home settings. LLMs may be able to determine whether a given piece of information is private, how that information should be processed, and with whom it may be shared. As privacy can be highly contextual (Nissenbaum, 2004, 2009), LLMs might help robots navigate the nuances of any given scenario.
Simultaneously, researchers in academia and industry are deploying LLMs and video language models (VLMs) on robots. These models equip robots with the tools to perceive and reason about their environments far beyond their previous capabilities. Researchers have demonstrated that LLMs can enhance robot capabilities by enabling natural language requests (Ahn et al., 2022), writing code to control robot movement on-the-fly (Liang et al., 2023), achieving interactive dialogue (Shinn et al., 2023), and improving navigation abilities (Hassan et al., 2024). Several companies have also introduced projects that utilize LLMs or VLMs to enable more intelligent robots, such as Google’s DeepMind-RT-2 (Brohan et al., 2023) and NVIDIA’s GR00T N1 (Bjorck et al., 2025). Additionally, researchers have evaluated LLM-powered robots through user studies to identify the potential benefits and weaknesses of these LLM frameworks (Yuan et al., 2024a; Murali et al., 2023; Kim et al., 2024a; Azeem et al., 2024). From these advancements, it is clear that LLMs and VLMs can offer significant benefits when integrated with robots.
@@
As LLMs are increasingly integrated into home robots, their ability to navigate nuanced privacy decisions regarding data collection, processing, and sharing becomes critical. Prior works have primarily focused on general LLM safety, such as LLMs’ vulnerability against jailbreak attacks (Zhuo et al., 2023). Few works, however, have investigated the privacy perceptions of LLMs and VLMs (Mireshghallah et al., 2023; Shao et al., 2024), and none, to our knowledge, have done so in the context of human-robot interaction (HRI). Given this gap, we investigate how LLMs help robots interpret privacy and make privacy decisions regarding data collection, processing, and sharing in home settings. Specifically, we raise the following research questions: RQ1: How does an individual’s privacy orientation influence their privacy expectations of robots in home settings?; RQ2: How well do state-of-the-art LLMs align with individuals’ privacy expectations?; and RQ3: How can we improve state-of-the-art LLMs’ conformity with individuals’ privacy expectations?
Through the development and evaluation of privacy scenarios with participants and LLMs, we show that state-of-the-art LLMs maintain a broad understanding of privacy sensitivity, but may not yet possess nuanced privacy awareness. In addressing our research questions, we present the following contributions:
  • We curate a dataset of 50 in-home scenarios grounded in contextual integrity (CI), and develop a survey to measure users’ privacy preferences of robots’ data collection and processing behaviors. Utilizing this dataset and survey, we conduct an online user study with 450 participants to evaluate users’ preferences of robot behaviors in home scenarios.
  • We assess the in-home privacy awareness of 10 state-of-the-art open-source and closed-source LLMs using four different prompting strategies. We find that while LLMs default to high privacy-preserving options, models with few-shot prompting are more capable of reflecting user preferences.
  • We discuss the design implications of LLM-powered robots, suggesting the potential of using LLMs as privacy controllers for robots. By promoting data collection indicators and adopting users’ privacy preferences collected through interaction, robots can enable context- and privacy-aware behavior tailored to an individual’s needs.
@@ 2. Related Works
As we aim to explore whether user privacy preferences align with LLM-powered robots, we discuss the following threads of existing literature: (1) conceptual approaches to privacy, including privacy orientations and contextual integrity, (2) privacy concerns related to robot data collection in homes, (3) the integration of LLM reasoning in robots, and (4) LLM privacy risks and benchmarking.
@@ 2.1. Conceptual Approaches to Privacy
Beyond this broad framework of privacy, many other works have attempted to assess the specific privacy attitudes, orientations, and personas of individuals as they engage with technology (Preibusch, 2013). These efforts have focused on online privacy (Elueze and Quan-Haase, 2018; Buchanan et al., 2007), out-of-device privacy (i.e., privacy violations through direct human observation) (Farzand et al., 2024), privacy perceptions broadly (Chignell et al., 2003; Giang et al., 2023), and the identification of privacy personas (Hrynenko and Cavallaro, 2024). Of particular note to this work is the Privacy Orientation Scale (POS) (Baruh and Cemalcılar, 2014), which captures an individual’s perceptions of privacy as a right, concerns about their own informational privacy, other-contingent privacy, and concerns about the privacy of others. These four subscales can then be used to classify a user’s privacy orientation as fitting one of three profiles, including privacy advocates, privacy individualists, and privacy indifferents. Other works have described the general challenge of defining privacy personas (Biselli et al., 2022). Given the substantial contextual dependence of privacy, domain-agnostic assessments of privacy perceptions may also struggle to generate consistent personas across domains.
@@ 2.2. Privacy in Robotics
While privacy has been explored in a myriad of fields, it is a growing area of concern in robotics (Denning et al., 2009; Guerrero, 2017; Rueben et al., 2018; Torresen, 2018; Pagallo et al., 2018; Lutz et al., 2019; Chatzimichali et al., 2020). Particular emphasis has been placed on robotics usage in homes (Levinson et al., 2024b, a; Fernandes et al., 2016; Urquhart et al., 2019), as these environments are highly private, contain substantial amounts of sensitive information, and offer inhabitants a sense of security. Given the capacity of some robots to capture data, autonomously navigate environments, and socially engage with humans, they pose a significant threat to privacy with unique challenges compared to many other technologies. Existing works have also focused on the inherent privacy-utility trade-off that exists when robots are introduced into new environments (Butler et al., 2015; Leite and Lehman, 2016). While users may have concerns about robot privacy violations, these may be tolerated if sufficient utility is provided by the robot (Lutz and Tamó-Larrieux, 2020).
Though these concerns are widespread, many prior works have proposed solutions that attempt to limit data collection, retention, and exposure (Sullivan and Mutlu, 2025). For example, low-resolution sensors (Kim et al., 2019) and purposeful sensor selection (Eick and Antón, 2020) have been proposed to limit unnecessary data collection, while sensor redirection (Shome et al., 2023) and privacy controllers (Tang et al., 2022) have been proposed to mitigate data exposure. While these efforts have uncovered new methods for preserving human privacy, the combination of LLMs and robots presents novel opportunities for privacy-preserving LLM decision making.
@@ 2.3. Robots & LLM Integration
While mainstream LLMs are a relatively recent development (Vaswani et al., 2017; Radford et al., 2018), LLM integration in robots has been a clear area of interest for HRI applications. Though this intersection of technologies is still in its infancy, several works have already utilized LLMs in robotics applications. These applications have been specialized, such as a chemistry assistant robot (Yoshikawa et al., 2023) and a health attendant robot (Kim et al., 2024b), as well as broad, such as a ‘robotic brain’ to manage memory and control (Mai et al., 2023). Other works have considered the potential of LLM-powered robots compared to other forms of LLMs (i.e., text- and voice-based) (Kim et al., 2024a). All of these works, however, only begin to explore the full potential of robot-LLM integration. Further study is needed to understand the implications of LLMs for perception, control, interaction, and, most relevant to this work, decision making (Zeng et al., 2023).
@@ 2.4. LLM Privacy Risks & Benchmarking
To assess the potential risks and benefits that LLMs present, researchers utilize benchmarking as a standardized evaluation method for LLM performance. Common evaluation domains include reasoning (Srivastava et al., 2022; Hendrycks et al., 2020; Yi et al., 2025), multilingualism (Hu et al., 2020; Liang et al., 2020), coding (Chen et al., 2021; Austin et al., 2021), and instruction following (Zheng et al., 2023; Dubois et al., 2025) as well as safety and privacy (Ruan et al., 2023; Yuan et al., 2024b; Zhang et al., 2023; Gu et al., 2024; Zhang et al., 2024b; Zharmagambetov et al., 2025; Zhu et al., 2024; Zhang et al., 2024a; Chang et al., 2025; Bagdasarian et al., 2024). PURE (Chi et al., 2023), PolicyQA (Ahmad et al., 2020), PrivacyQA (Ravichander et al., 2019), and GenAIPABench (Hamid et al., 2023) have been performed to evaluate privacy-policy comprehension using dedicated datasets. Of particular relevance to this work, Mireshghallah et al. (2023) first raised the question of whether LLMs can reason about CI. These authors utilized their benchmark, ConfAlde, to show that GPT-4 violates CI norms in 39% of evaluated scenarios where humans would withhold information. Most recently, PrivacyLens (Shao et al., 2024) introduced vignette-based questions to survey LLMs’ privacy-norm awareness. However, existing works study whether an LLM can decide on the appropriateness of sharing data using tools from CI. In this work, we aim to explore how these models make decisions, and how well their responses align with human preferences in household HRI contexts.
@@
Robot Involvement Robot enters part-way, Robot present from the beginning, Robot collects data from afar (i.e., spying and eavesdropping)
@@
Table 1. Dimensions of Contextual Privacy—Dimensions of privacy relevant to household HRI. These dimensions were derived from literature on Contextual Integrity and expanded to include subdimensions and factors that are relevant to HRI.
@@ 3. Developing Privacy Scenarios
To address our research questions, we develop a set of privacy-relevant scenarios that allow participants to report their level of concern with a robot collecting user data and their preferences for how that robot should behave following data collection. Here, we describe our process to develop these scenarios, including the use of two online user studies to crowdsource specific scenario characteristics and robot behaviors. The studies were approved by our Institutional Review Board, and all participants provided informed consent prior to their participation.
@@ 3.1. Dimensions of Contextual Integrity
We first reviewed existing literature on general in-home privacy concerns (Choe et al., 2011) and Contextual Integrity (Nissenbaum, 2004). From this literature, we became oriented with a comprehensive list of the many in-home activities individuals wish to keep private. Additionally, we distilled the primary dimensions of contextual privacy, including context, actors, attributes, transmission principles, and purposes. From each of these dimensions, we further defined subdimensions relevant to privacy in HRI. These expanded dimensions include type of location, robot involvement, number of users, age of users, relation of users, medium of data, protected status of data, information flow, and task dependence. Table 1 presents a full list of all original dimensions, expanded dimensions, and factors.
While most of these dimensions and their factors are fairly intuitive, the factors of protected status and medium of data are worth discussing further. Our original intent behind protected status was to capture the sensitivity of the information involved in a given scenario. Given how subjective sensitivity may be, we chose to instead define protected status as either protected or non-protected based upon legal norms. For example, information we consider to be protected includes health data (for Civil Rights, 2003), academic records (U.S. Congress, 1974), and legal records (United States Courts, 2008), among others (Internal Revenue Service, 1976; U.S. Congress, 1999). Conversely, we consider information such as vacation plans or intimacy between partners to be non-protected. Additionally, we considered only specific mediums of data that we believed an LLM or VLM could reliably recognize. These data types include visuals of the human body in compromising situations (e.g., bathing, toileting, or intimacy), visuals of device screens, audio of conversations, and visuals of physical documents. These broad categories were selected due to the wide variety of information each of them can convey.